RC meeting and Run Plan, Tuesday Oct. 3, 2023
Return to the RC Meeting Calendar
- 1 RC Meeting Information
- 2 Useful Links
- 3 Procedures
- 4 Status Update
RC Meeting Information
Time: Monday - Friday at 3:30pm
- Location: Counting house conference room
- ZoomGov Meeting Connection Information
Hall A call-in instructions:
- Call RC first.
- Techs on site 7am-4:30pm Mon-Thu, 7-11am Fri. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
- After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.
Hall Access procedure:
- Access check list (binder on the counting room desk) needs to be filled out before making an access.
- Scan (take pictures with phone) and post it to the logbook.
- Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.
New instruction for shift crew when changing target:
- Beam restoration instructions have been updated. Printed version is in the counting room.
- When we change target, make sure the SOFs of Target BPMs are set to the correct values.
- SOFs for different targets: Look in the 'HOW TOs' page on the SBS Wiki for details (Note: SOFs for major target positions are on the Control Room whiteboard.)
After the access was completed yesterday, the beam was not restored until mid-way through SWING shift. Due to fighting with the DAQ, production was not ready to restart until 11:35 pm. During OWL shift, production data was taken at 40uA to let the target polarization continue to recover from the 8 hours that the lasers were off. DAY shift increased the beam current to 45uA, where it has stayed since.
- The 4-pass separator work took 12 hours to complete, so beam recovery couldn't start until 9:15pm.
- Since then, the beam has been stable at 45uA.
- Hall B plan on restarting during SWING shift, but this isn't set in stone.
- No maintenance day on the MCC white board (until 10/10/2023).
Data collected :
- Accumulated corrected charge collected: 0.972C from midnight until 1:30pm.
- 0.685C during OWL
- 0.287C during DAY (until 1:30pm)
Hall A Subsystems Status
- The target span up overnight, and is stable.
- Latest polarization measurement: 38.3% (https://logbooks.jlab.org/entry/4191731).
Hall access requests
- If anyone wants to make an access, let me know! There is no planned maintenance day before Tuesday 10th, so any accesses need to be opportunistic.
- Still need to fill open slots
- 10/5 - need a TO for OWL
- Please check when your shifts are!
Please sign up for any open shifts!
Short Term Run plan
- Continue running at 45uA.
- If the target polarization drops significantly, lower the current to 40uA.
==> List of items in case He3 production running isn’t ready:
- Take data on reference cell at 10 atm H2 (portions that add up to 1 day) [85% (4200 elastic events) more data required for Bigbite momentum calibration]
- BigBite Sieve slit => only during day (Techs needed)
- with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
- 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
- BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)